OpenStudy (anonymous):

what is the measure of an interior angle of a regular polygon is 170°. How many sides make up the regular polygon

OpenStudy (anonymous):

what is the meaning of the first part?

OpenStudy (anonymous):

can you rewrite it please ...

OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 170°. How many sides make up the regular polygon?

OpenStudy (blockcolder):

Each interior angle of a regular n-gon is of measure (180(n-2))/n. Let this equal to 170, then find n.

OpenStudy (anonymous):

Can you dumb it down a little for me its like 3 o'clock in the morning over here so im confused lol

OpenStudy (blockcolder):

\[\frac{180(n-2)}{n}=170\\ 180n-360=170n\\ 180n-170n=360\\ 10n=360\\ n=36\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

If the interior is 170 then the exterior is 10 and sides = 360/10 = 36